A local gasoline dealership in a small town wants to estimate the average amount of gasoline that people in that town use in a 1-week period. The dealer asked 44 randomly selected customers to keep a diary of their gasoline usage, and this information produced the following data on gas used (in gallons) by these people during a 1-week period.

23.1 269 14.5 105 136 229 96 31 25.8 18.3 121 52 100 23.5 18.0 05 76 21.6 205 83 18.9 155 142 10.5 266 23.5 71 74 138 11.8 13Construct a 95% confidence interval for the average weekly gas usage by people in this town. Use the t distribution.
Answer #1

C4 fx =STDEV(A1:A44) A B D E F 1 2 3 14.48409 6.484039 4 5 6 7 8 23.1 13.6 25.8 10 7.6 18.9 26.6 23.8 12.3 9 10 15.8 11 21 12Mean= 14.484 Standard deviation=6.484. CI= Xt 10.01) S Et0.05.45) 6.484 = 14.484 (2.017) = 14.484 +1.972 = (14.484 -1.972,14.The 95% confidence interval for the average weekly gas usage by people in this town is 12.512<u<16.456.

answered by: ANURANJAN SARSAM